# Reinforcement Learning with Action Chunking

### v1

@@ 2 Related Work
Action chunking is a technique popularized by roboticists for imitation learning (IL), where the policy predicts and executes a sequence of actions in an open-loop manner (“an action chunk”) [82]. Action chunking has been shown to improve policy robustness [82, 22, 8], and handle non-Markovian behavior in offline data [82]. Existing RL methods that incorporate action chunking typically focus on fine-tuning a policy pre-trained with imitation learning [57, 60]. Tian et al. [71] propose to learn a critic on action chunks by integrating $n$-step returns with a transformer. However, their method only applies chunking to the critic, while still optimizing a single-step actor.
@@ 4.1 Q-learning on a temporally extended action space
The first design principle of Q-chunking is to apply $Q$-learning on the temporally extended action space. Unlike normal 1-step TD-based actor-critic methods, which train a Q-function $Q(s_{t},a_{t})$ and a policy $\pi(a_{t}|s_{t})$, we instead train both the critic and the actor with a span of $h$ consecutive actions:  [^1] We use $\mathbf{a}_{t:t+h}$ to denote a concatenation of $h$ consecutive actions: $\begin{bmatrix}a_{t}&a_{t+1}&\cdots&a_{t+h-1}\end{bmatrix}\in\mathbb{R}^{Ah}$ for notation convenience. This is similar for $\mathbf{s}_{t:t+h}$ and $\mathbf{r}_{t:t+h}$.
@@ 4.3 Practical implementations
While it is possible include the KL as part of the loss, estimating the KL divergence or log probability for flow models is practically challenging. Instead, we use best-of-$N$ sampling [67] to maximize $Q$-value while imposing this KL constraint implicitly altogether. Practically, this involves first sampling $N$ action chunks from the learned behavior policy $f_{\xi}(\cdot|s_{t})$,
@@
While QC is simple and easy to implement, it does come with some additional computational costs (sampling $N\times$).
We present a variant of our method below that leverages a cheaper off-the-shelf offline/offline-to-online RL method, FQL [54].

### v2

@@ 2 Related Work
Action chunking is a technique popularized by roboticists for imitation learning (IL), where the policy predicts and executes a sequence of actions in an open-loop manner (“an action chunk”) [87]. Action chunking has been shown to improve policy robustness [87, 22, 8], and handle non-Markovian behavior in offline data [87]. Existing RL methods that incorporate action chunking typically focus on fine-tuning a policy pre-trained with imitation learning [59]. Tian et al. [75] propose to learn a critic on action chunks by integrating $n$-step returns with a transformer. However, their method only applies chunking to the critic, while still optimizing a single-step actor. Li et al. [37] also observe that learning a critic over short action chunks removes the off-policy bias in $n$-step return backups, leading to more stable and effective value learning. There are key differences between their work and ours — Li et al. [37] operate in the online episodic RL setting [81, 30] and use Gaussian policies to predict parameters of motion primitives (MP) [61, 52], which are then used to generate full action sequences at the beginning of each episode. In contrast, we operate in the conventional offline-to-online RL setting and leverage more expressive flow-matching-based policies (as we find Gaussian policies to be ineffective as shown in Figure 2) to predict short action sequences directly in the raw action space. Seo and Abbeel [63] also train critics on action chunks and impose a behavior cloning loss that aligns with the principles behind Q-chunking. The key distinction from our work lies in their use of a multi-level, factorized critic architecture [64], which generates and refines action chunks from coarse to fine granularity via iterative discretization. At each level, the action space is discretized into bins, and the Q-function is modeled independently for each action dimension and timestep, conditioned on the whole action chunks predicted at the previous coarser level. While this factorized critic design enables tractable value-maximizing action sampling, it imposes strong structural assumptions on the action space, limiting policy expressiveness at each refinement level. In contrast, we make no such assumptions in our recipe which allows us to derive two general algorithms where both the critic and policy operate directly on action chunks without requiring factorization or iterative discretization/refinement.
@@ 4.1 Q-learning on a temporally extended action space
The first design principle of Q-chunking is to apply $Q$-learning on the temporally extended action space. Unlike normal 1-step TD-based actor-critic methods, which train a Q-function $Q(s_{t},a_{t})$ and a policy $\pi(a_{t}|s_{t})$, we instead train both the critic and the actor with a span of $h$ consecutive actions:  [^1] We use $\mathbf{a}_{t:t+h}$ (or simply $\mathbf{a}_{t}$) to denote a concatenation of $h$ consecutive actions: $\begin{bmatrix}a_{t}&\cdots&a_{t+h-1}\end{bmatrix}\in\mathbb{R}^{Ah}$ for notation convenience. This is similar for $\mathbf{s}_{t:t+h}$ and $\mathbf{r}_{t:t+h}$.
@@ 4.3 Practical implementations
While it is possible include the KL as part of the loss, estimating the KL divergence or log probability for flow models is practically challenging. Instead, we use best-of-$N$ sampling [71] to maximize $Q$-value while imposing this KL constraint implicitly altogether. Practically, this involves first sampling $N$ action chunks from the learned behavior policy $f_{\xi}(\cdot|s)$,
@@
While QC is simple and easy to implement, it does come with some additional computational costs (sampling $N\times$). We present a variant of our method below that leverages a cheaper off-the-shelf offline/offline-to-online RL method, FQL [56].